OpenStudy (anonymous):

6x-18y=-3 -2x+6y=-12 solve using either substitution or elimination

OpenStudy (kira_yamato):

6x-18y=-3 -------(1) -2x+6y=-12 ------(2) From (2): x = 3y + 6 --------(3) Plug (3) into (1): 6(3y + 6) - 18y = -3 18y + 36 - 18y = -3 36 = -3 (No solution)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

6x-18y=-3 -------(1) -2x+6y=-12 ------(2) From (2): x = 3y + 6 --------(3) Plug (3) into (1): 6(3y + 6) - 18y = -3 18y + 36 - 18y = -3 36 = -3 (No solution)

OpenStudy (anonymous):

thanks